total no of matrices can be  formed by 6 diff numbers

Grade:12

3 Answers

Chunduru Vaishnav
6 years ago
  1. Number of 1×1 matrix formed by 6 digits are 6
  2. Number of 2×2 matrix formed by 6 digits are 6​4 becase 2×2 matrix has 4 elements since repetition is allowed each element can take any of the six number
  3. Similarly for 3×3 matrix is 69
And total number of matrives are 6+64+69+------.       = Infinity
diwakar raturi
6 years ago
no. of matrix of 1*1=6no.of of matrixof 2*2=6*6*6*6=6^4 so similarlytotal no of matrix=6+6^4+6^9+`.....…= infinity
om
4 years ago
possible order will be 1*6,6*1,2*3 and 3*2
now for each order we can arrange element in form of factorial 6
thus total number of matrix wiil be 4 *factorial 6
note;*multiplication